Evaluate without using math table

Sqrt(Rationalize(0.81)*10^(-5)/(Rationalize(2.25)*10^7))



1
Expert's answer
2021-11-08T20:28:24-0500
"\sqrt{\dfrac{0.81\cdot10^{-5}}{2.25\cdot10^7}}=\dfrac{0.9}{1.5}\cdot10^{-6}=0.6\cdot10^{-6}"

"=6\cdot10^{-7}"